OpenStudy (lgbasallote):

factor out: \[4y^2 + 16x^2 + 24xz^2 - 8xy + 9z^4 - 12yz^2\]

OpenStudy (anonymous):

factor out what?

OpenStudy (anonymous):

are you sure its 8xy?

OpenStudy (anonymous):

Dr. House? your wanted at your question...

OpenStudy (anonymous):

\[\large 1*(4y^2 + 16x^2 + 24xz^2 - 8xy + 9z^4 - 12yz^2)\]

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

that's the expression..no mistakes

OpenStudy (anonymous):

see \[a^2 + b^2 +c^2+2ab+2bc+2ca = (a+b+c)^2\] In your question a = 4x b = -2y c = 3z :)

OpenStudy (anonymous):

but then it should be 16xy

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So add and subtract 8xy

OpenStudy (lgbasallote):

hmm i think my teacher mentioned a correction about this -_- i cant recall that was one year ago lol sorry :P

sam (.sam.):

lol

OpenStudy (anonymous):

So it should be \[(4x-2y+3z^2)^2 +8xy\]